OpenStudy (anonymous):

. Describe the graphs of the functions f(x) = 2^x – 3 and g(x) = –2^x – 3. Compare and contrast the domain and range of f(x) and g(x). now problem!

OpenStudy (anonymous):

i know that f(x)= 2^x-3 is going up and g(x)=-2^x -3 is going down.

OpenStudy (anonymous):

what about the domain and rang?

OpenStudy (anonymous):

they're both negative now?

OpenStudy (anonymous):

 like that?